pointA(-1,y) and B(5,7) lie on the circle with centre O(2,-3y).Find the values of y.Hence findd the radius of the circle

Grade:12

2 Answers

Ravi
9 years ago
OA=OB → Find the value of y from this eqn using distance formula for both OA and OB. After solving for y, Find either OA or OB to be the radius of the circle.
Deepak Jindal
7 years ago
Given, Centre of the circle is O (2, 3).A (4, 3) and B (x, 5) are two points on the circle.∴ OA = OB (Radius of the circle)Squaring on both sides, we get∴ 4 + 0 = (x – 2)2 + 4⇒ (x – 2)2 = 0⇒ x – 2 = 0⇒ x = 2Thus, the value of x is 2.